A solution of malonic acid, H2C3H2O4, was standardized by titration with 0.1000 M NaOH solution. If 20.76 mL of the NaOH solution is required to neutralize completely 12.95 mL of the malonic acid solution, what is the molarity of the malonic acid solution ?

H2C3H2O4+ 2NaOH > Na2C3H2O4 + 2H2O
1
Expert's answer
2019-12-11T07:54:23-0500

  Xmol        0.002076mol

H2C3H2O4 + 2NaOH> Na2C3H2O4 + 2H2O

   1mol            2mol


n(NaOH) = C(NaOH) × V(NaOH) = 0.1000M × 0.02076L = 0.002076mol.


n(H2C3H2O4) = X = 0.002076 × 1 / 2 mol = 0.001038mol.


C(H2C3H2O4) = n(H2C3H2O4)/ V(H2C3H2O4) = 0.001038mol/0.01295L = 0.080154M ≈ 0.0802M.